OpenStudy (caitlinnr14):

WILL GIV MEDAL What is the slope of the line 5x - 2y = -10?

OpenStudy (caitlinnr14):

CONFUSED

sam (.sam.):

In the form Y=MX+B, M is the slope, so could you find a way to rearrange 5x - 2y = -10 to that form?

OpenStudy (caitlinnr14):

?

OpenStudy (caitlinnr14):

im so confused im sorry

OpenStudy (littlebird):

Get the y by itself on one side

OpenStudy (uri):

@caitlinnr14 Formula is y=mx+b So first arrange your eq.

sam (.sam.):

Like this, \[5x - 2y = -10\] Add \(\color{blue}{+2y}\) to both sides and add \(\color{red}{+10}\) to both sides \[5x\color{blue}{-2y+2y}\color{red}{+10}=-10\color{red}{+10}\color{blue}{+2y}\] You get \[5x+10=2y\] Same as \[2y=5x+10\] Divide 2 both sides and you'll get a similar form of y=mx+b

OpenStudy (uri):

5x - 2y = -10? -2y=-5x-10

OpenStudy (caitlinnr14):

wait is it -2y=-5x-10 or 2y=5x+10 @uri @.Sam.

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

5x - 2y = -10? -2y=-5x-10

OpenStudy (caitlinnr14):

Which one? @jonjenkins7653

OpenStudy (caitlinnr14):

oh ok

OpenStudy (uri):

What i said @caitlinnr14

OpenStudy (anonymous):

yeah thats right

OpenStudy (uri):

Both are correct @caitlinnr14

OpenStudy (caitlinnr14):

so the slope would be -2/5? i dont know if thts rightttt

sam (.sam.):

Both 2y=5x+10 and -2y=-5x-10 is correct if you're not dealing with inequalities

OpenStudy (anonymous):

2y=5x+10 and -2y=-5x-10 is correct

OpenStudy (caitlinnr14):

oh ok @.Sam.

OpenStudy (littlebird):

A negative divided by a negative makes a positive

OpenStudy (anonymous):

those are both correct

OpenStudy (caitlinnr14):

oh sry thx i made a mistake @Littlebird

sam (.sam.):

Divide 2 to both sides \[y=\frac{5}{2}x+5\] \[y=mx+b\] m is slope, so what is slope?

OpenStudy (caitlinnr14):

5?

sam (.sam.):

Careful, \[y=(\frac{5}{2})x+5\]

OpenStudy (caitlinnr14):

ohhhhhh okkk

OpenStudy (caitlinnr14):

I got it noe thank yo so much i guess i get confused at the part where you convert it to y=mx+b

sam (.sam.):

Oh its simple its just direct conversion :)

OpenStudy (caitlinnr14):

thank you

sam (.sam.):

yw :)